# A Survey on Intent-aware Recommender Systems

### v1

@@ Abstract.
Recommender Systems
copyright: acmcopyrightjournalyear: 2024journal: TORSjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Information systems Recommender systems
@@ 1. Introduction
Given these diverse sets of potential short-term intents, a recommendation algorithm that considers all of our past preferences independently of the user’s current situation and goals may thus not be optimal. Since it can be difficult to correctly guess a user’s specific intent when accessing the service, a common solution in practice is to provide users with several recommendation lists on the landing page of the service, each of them addressing a different potential intent. Multiple recommendation lists are commonly used by media streaming providers such as Netflix or Spotify and on e-commerce sites like Amazon.com (Gomez-Uribe and Hunt, 2015; Mehrotra et al., 2019; Loepp, 2023). Alternatively, some services try to diversify their single-list recommendations to cover multiple potential user intents. They may, for example, both comprise both items that are assumed to help user discover new things but also deliberately include a small number of items in the recommendations that the user has recently viewed, e.g., to allow users to conveniently continue their previous shopping session (Lerche et al., 2016).
@@
In the academic literature, the problem of diverse and time-varying user needs has been identified and addressed in different ways over time. In the realm of the traditional way of building recommender systems based on a given user-item interaction matrix, common approaches include the intentional diversification of the recommendations, the consideration of temporal information and interest drifts, or the utilization of contextual information like the user’s current location and time (Campos et al., 2014; Bogina et al., 2022; Adomavicius and Tuzhilin, 2015; Kaminskas and Bridge, 2016). More recently, academic research has largely shifted from the traditional “matrix completion” recommendation setting to sequence-aware and session-based recommendation problems (Quadrana et al., 2018; Wang et al., 2021a). In these approaches, the recommendations that are generated are commonly determined by considering the most recent interactions that are observed for a given user. The main underlying assumption of these approaches is that these recent interactions are a good reflection the user’s underlying short-term intents.
@@
However, while the most recently observed user actions may be indicative of the underlying user intents, researchers have recently tried to incorporate intent-awareness into recommender systems in a more explicit manner. Importantly, reports from industrial applications emphasize the strong potential of adapting recommendations to the predicted intent(s) of the users. For example, a field test in the e-commerce domain (Shi and Ghedira, 2017) revealed that relevant business KPIs (key performance indicators) like purchase rates can be markedly improved when the recommendation system dynamically switches its strategy depending on the currently assumed user intent. Likewise, a study in the music streaming domain by Spotify (Mehrotra et al., 2019) investigated the connections between observed user interactions, underlying intents and user satisfaction. One main conclusion of this work is that understanding user intents can be crucial to predict user satisfaction, and that different interaction signals are relevant as predictors, depending on the user’s underlying intent.
@@
Given the high practical relevance of such approaches, the aim of our present work is to review the growing body of literature on intent-aware recommender systems (IARS). For this purpose, we have identified a larger set of recent works based on a semi-systematic literature search, and we have categorized these works along different dimensions. A particular focus of the survey lies on the technical approaches to incorporate intent-awareness into the system, such as diversification techniques, intent prediction models or latent intent modeling approaches. We furthermore review common application domains and evaluation procedures for IARS, which revealed a very strong focus on offline experimentation with an emphasis on prediction accuracy as the main metric. Finally, our survey helped us to identify open gaps in the literature and future research directions. Specifically, the increasing availability of additional pieces of information about users and their behavior, about item features and contextual factors opens continuously new opportunities to better predict a user’s specific intent when interacting with the service. Furthermore, we see significant potential in the explicit consideration of application-specific intents within future IARS.
@@
The paper is organized as follows. Next, in Section 2, we review different notions of intent-awareness as found in the literature, and we delineate our work from related streams of research. Furthermore, we describe our methodology of identifying relevant papers and discuss the relationship of our work to other surveys. In Section 3, we then categorize and discuss different technical approaches to build intent-aware recommender systems. Section 4 subsequently provides a landscape of current research in terms of the used types of data, evaluation approaches, and application domains. The paper ends with a discussion of research gaps and future research directions.
@@ 2. Terminology and Research Methodology
In this section, we first elaborate on terminological questions and provide a definition of intent-aware recommender systems. Then, we explain our research methodology and discuss previous works.
@@ 2.1. Terminology and Definitions
According to the Cambridge Dictionary [^1] https://dictionary.cambridge.org/, the noun ‘intent’ refers to “the fact that you want and plan to do something” or “the intention to do something”. A similar action-oriented definition is provided by Wikipedia [^2] https://en.wikipedia.org/wiki/Intent_(disambiguation), where intent is described as “an agent’s specific purpose in performing an action or series of actions.” Common synonyms include the terms ‘purpose’, ‘plan’, ‘goal’, or ‘aim’. We note that the word ‘intention’ is often used synonymously with ‘intent’ as well, even though the two terms are not entirely interchangeable, with the term ‘intent’ suggesting a “clear[er] reasoning or great deliberateness.” [^3] https://www.britannica.com/dictionary/eb/qa/intent-and-intention Intuitively, an intent-aware RS should therefore take into account why a user is interacting with a system or which goal she or he wants to achieve, thus, complementing the traditional approach of predicting which items might be relevant for the specific user.
@@
In the surveyed literature researchers rarely specify explicitly what they mean by the terms ‘intent’ or ‘intent-awareness’ or refer to humanities and social sciences literature, e.g., (Wang et al., 2020b, 2022e). More frequently, they provide intuitive examples to illustrate their notion of an intent-aware system. Different interpretations of intent-awareness can be found, and these interpretations are typically tied to the technical approaches that are presented in the respective papers to implement intent-awareness in the system; we will discuss these technical approaches in more detail in Section 3. Following the spirit of early diversity-oriented IARS approaches, the authors of (Kaya and Bridge, 2019a) for example state that “[IARS] ensure that the set of recommendations contains items that cover each of the user’s interests […].” More recent works like (Wang et al., 2020c) relate intents with observed actions and provide examples, in this case for shopping intents in the e-commerce domain, e.g., “passing time” or “shopping for others”. In a number of other works, e.g., (Chen et al., 2020, 2019), the term ‘intent’ is mostly considered to be equivalent with (observed) interest, e.g., in specific topics or item categories. Finally, in some works like (Jannach et al., 2017), the term (short-term) ‘intent’ is merely used to describe a sequential recommendation approach that puts more weight on the last observed user interactions. Thus, intent is sometimes only equated with observed user actions, without additional consideration about the users’ underlying motivations for these actions.
@@
Given the examples from the previous paragraph, we can roughly categorize existing IARS approaches into three classes. First, a comparably small number of the surveyed works are based on a small set of pre-defined and application-specific explicit user intents, e.g., “discover new music to listen now” or “find background music to listen to” (Mehrotra et al., 2019). In such an approach, one can for example apply different recommendation strategies depending on the currently assumed intent. Second, some other works provide examples of such explicit intents for illustration, e.g., “shopping for others” (Wang et al., 2020c), but are actually based on the concept of ‘latent’ intents. In fact, much of the surveyed literature is based on the idea of incorporating (additional) latent variables in the models to be able to account for a typically small set of possible user intents. Third, there are a number of approaches, as mentioned above, in which intent is considered to be roughly equivalent to observed past interests in categories, certain item features, or even individual items. In such approaches, additional variables are commonly introduced, which are however not assumed to directly correspond to specific underlying intents in the sense of goals a user wants to achieve. Instead, such variables may, for instance, model the user’s time-varying interest in certain item categories or other assumed relationships between individual items. We denote such approaches implicit intent models.
@@
An IARS is a recommender system that aims at capturing, at least implicitly, the users’ underlying current motivations and goals in order to support them.
Our definition is purposely rather broad, yet it expresses that the design of any IARS should be at least conceptually driven by the ambition to support the users’ current motivations. Given the breadth of different notions of intent-awareness mentioned above, a multitude of technical ways exist to implement intent-awareness. On one end of the spectrum, we for example have approaches that are based on an explicit list of pre-defined application-specific intents. In contrast, there are also works in which any sequential recommender system that puts more emphasis on the most recently observed interactions can be seen as a mechanism to implement an (implicitly) intent-aware system, e.g. (Jannach et al., 2017).
@@ 2.2. Research Methodology
We adopted a semi-systematic approach to identify relevant papers. First, we queried services like Google Scholar and digital libraries for scientific papers that mention the terms ‘intent’ or ‘intention’ together with the term ‘recommend’ in the title or abstracts. We then screened these papers for relevant ones and followed the references mentioned in these papers to identify additional papers. We only retained papers that had undergone a peer-review process, i.e., we excluded preprints from our analysis. Furthermore, a paper was considered relevant if its contribution was motivated by the idea of intent-awareness in the sense of our definition from above even though it did not explicitly contain our search terms in the title or abstract. As such, a small number of papers was considered that emphasized on identifying and addressing the user’s current aims and interests.
We also identified a number of works which mention the term ‘intent’ in the title or abstract, but not in a context that is relevant for our survey. For example, some works aim at predicting if a consumer will make a purchase in an ongoing session or not, e.g., (Requena et al., 2020). In such works, the goal is not to make recommendations that match a certain intent, as done in IARS, but to assess the general propensity of a consumer to make a purchase, see also (Koniew, 2020). In some other works, the term ‘intent’ is used in approaches where behavioral logs are used to estimate the propensity of a consumer to purchase a certain item or an item from a certain category (He et al., 2014a; Bhattacharya et al., 2017; Sulikowski et al., 2021). Again, such approaches do not match our definition of IARS, which is relating to a consumer’s underlying goals and motivations.
Overall, we ended up with 81 papers that we considered relevant and which we categorize later in Section 3.555The list of considered papers can be found online at https://docs.google.com/spreadsheets/d/1ccEDpdJ2MXcsEUCJHWuamqUjDeuwB_b1ttxdstKhNw8/. Finally, our literature search also surfaced surveys that did not propose a technical contribution or reported an experimental study. We discuss these papers separately next.
@@
Surveys on related the areas, as discussed above, such as time-aware, contextual, and sequential recommender systems, can be found in (Campos et al., 2014; Bogina et al., 2022; Adomavicius and Tuzhilin, 2015; Kaminskas and Bridge, 2016; Quadrana et al., 2018; Wang et al., 2021a). Early intent-aware recommendation systems were inspired by existing approaches to search result diversification in the field of Information Retrieval, see (Santos et al., 2015) for a related survey. Indeed, considering the users’ intents, e.g., in terms of their ‘informational needs’ has a long history in Information Retrieval, see (Broder, 2002) for an early taxonomy. A review of intent-awareness in the particular context of multimedia information retrieval can be found in (Kofler et al., 2017). Like in our survey, Kofler at al. (Kofler et al., 2017) found that various notions of intent-awareness exist in the literature. Our definition of intent-awareness is similar to theirs by focusing on the ‘why’ dimension, i.e., the reason and goal behind a search, and not the ‘what’ dimension of information needs. In the area music information retrieval, Schedl et al. (Schedl et al., 2018) reviewed current challenges and visions in music recommendation research. While their review work is not primarily focused on the topic of user intents, they identify an improved understanding of listening intents and a user’s purpose as a grand challenge in this area. In that context, they also mention that creators may have specific intents in mind when designing a playlist (e.g., for relaxing).
Probably most similar to our survey is the work by Rafique et al. (Rafique et al., 2023), which also focuses on intent-aware recommender systems. They, however, study the current and future role of IARS from the lens of smart cities. A particular emphasis is put on the role of recent technological developments in this context, e.g., the increasing usage of wearable and intelligent devices, the Internet of Things or Edge Computing approaches. As a result, new opportunities arise to build future intent-aware systems based, for example, on real-time activity, emotion tracking or the explicit elicitation of user intents based on new sensors and devices. In their survey, the authors thus particularly focus on case studies of IARS in smart cities and on key requirements in this context.
@@ 3. Technical Approaches for IARS
In this section, we categorize the considered studies on intent-aware recommender systems according to their underlying technical approaches. We categorize the underlying approaches as illustrated in Figure 1.
@@
We note that some papers may fall into more than one category, e.g., latent-intent modeling approach with the goal of diversification. In such cases, we categorized these works according to what we identified as their main conceptual contribution.

### v3

@@ Abstract.
Recommender Systems, Intent-Awareness, Survey
copyright: acmcopyrightjournalyear: 2024journal: TORSjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3700890ccs: Information systems Recommender systems
@@ 1. Introduction
Given these diverse sets of potential short-term intents, a recommendation algorithm that considers all of our past preferences independently of the user’s current situation and goals may thus not be optimal. Since it can be difficult to correctly guess a user’s specific intent when accessing the service, a common solution in practice is to provide users with several recommendation lists on the landing page of the service, each of them addressing a different potential intent. Multiple recommendation lists are commonly used by media streaming providers such as Netflix or Spotify and on e-commerce sites like Amazon.com (Gomez-Uribe and Hunt, 2015; Mehrotra et al., 2019; Loepp, 2023). Alternatively, some services try to diversify their single-list recommendations to cover multiple potential user intents. They may, for instance, comprise items that presumably constitute novel and serendipitous discoveries to a given user while, at the same time, a small number of items the user has recently already seen is also included in recommendation lists. Due to the latter the user could conveniently continue a previous shopping session (Lerche et al., 2016).
@@
In the academic literature, the problem of diverse and time-varying user needs has been identified and addressed in different ways over time. In the realm of the traditional way of building recommender systems based on a given user-item interaction matrix, common approaches include intentional diversification of recommendations, temporal information and interest drifts, or utilization of contextual information like the user’s current location and time (Campos et al., 2014; Bogina et al., 2022; Adomavicius and Tuzhilin, 2015; Kaminskas and Bridge, 2016). More recently, academic research has largely shifted from the traditional “matrix completion” recommendation setting to sequence-aware and session-based recommendation problems (Quadrana et al., 2018; Wang et al., 2021a). In these approaches, generated recommendations are commonly determined by considering the most recent interactions observed for a given user. The main underlying assumption of these approaches is that these recent interactions are a good reflection the user’s underlying short-term intents.
@@
However, while the most recently observed user actions may be indicative of the underlying user intents, researchers have recently tried to incorporate intent-awareness into recommender systems in a more explicit manner. Importantly, reports from industrial applications emphasize the strong potential of adapting recommendations to the predicted intent(s) of users. For instance, a field test in the e-commerce domain (Shi and Ghedira, 2017) revealed that relevant business KPIs (key performance indicators) like purchase rates can be markedly improved when the recommendation system dynamically switches its strategy depending on the currently assumed user intent. Likewise, a study in the music streaming domain by Spotify (Mehrotra et al., 2019) investigated the connections between observed user interactions, underlying intents and user satisfaction. One main conclusion of this work is that understanding user intents can be crucial to predict user satisfaction, and that different interaction signals are relevant as predictors, depending on the user’s underlying intent.
@@
Given the high practical relevance of such approaches, the aim of our present work is to review the growing body of literature on intent-aware recommender systems (IARS). For this purpose, we have identified a larger set of recent works based on a semi-systematic literature search, and categorized these works along different dimensions. A particular focus of this survey lies on the technical approaches to incorporate intent-awareness into the system, such as diversification techniques, intent prediction models or latent intent modeling approaches. We furthermore review common application domains and evaluation procedures for IARS, which revealed a very strong focus on offline experimentation with an emphasis on prediction accuracy as the main metric. Finally, our survey helped us to identify open gaps in the literature and future research directions. Specifically, the increasing availability of additional pieces of information about users and their behavior as well as on item features and contextual factors opens continuously new opportunities to better predict a user’s specific intent when interacting with the service. Furthermore, we see significant potential in the explicit consideration of application-specific intents within future IARS.
@@
The paper is organized as follows. Next, in Section 2, we review different notions of intent-awareness as found in the literature, and we delineate our work from related streams of research. Furthermore, we describe our methodology of identifying relevant papers and discuss the relationship of our work to other surveys. In Section 3, we then categorize and discuss different technical approaches to build intent-aware recommender systems. Section 4 subsequently provides a landscape of current research in terms of the used types of data, evaluation approaches, and application domains. The paper finishes with a discussion of research gaps and future research directions.
@@ 2. Terminology and Research Methodology
In this section, we first elaborate a definition of intent-aware recommender systems based on how they are actually reported in the literature. Then, we explain our research methodology and discuss previous works.
@@ 2.1. Terminology and Definitions
According to the Cambridge Dictionary [^1] https://dictionary.cambridge.org/, the noun ‘intent’ refers to “the fact that you want and plan to do something” or “the intention to do something”. A similar action-oriented definition is provided by Wikipedia [^2] https://en.wikipedia.org/wiki/Intent_(disambiguation), where intent is described as “an agent’s specific purpose in performing an action or series of actions.” Common synonyms include the terms ‘purpose’, ‘plan’, ‘goal’, or ‘aim’. We note that the word ‘intention’ is often used synonymous with ‘intent’ as well, even though the two terms are not entirely interchangeable, with the term ‘intent’ suggesting a “clear[er] reasoning or great deliberateness.” [^3] https://www.britannica.com/dictionary/eb/qa/intent-and-intention Intuitively, an intent-aware RS should therefore take into account why a user is interacting with a system or which goal she or he wants to achieve, thus, complementing the traditional approach of predicting which items might be relevant for the specific user.
@@
In the surveyed literature researchers rarely specify explicitly what they mean by the terms ‘intent’ or ‘intent-awareness’ or refer to humanities and social sciences literature, e.g., (Wang et al., 2020b, 2022e). More frequently, they provide intuitive examples to illustrate their notion of an intent-aware system. Different interpretations of intent-awareness can be found, and these interpretations are typically tied to the technical approaches that are presented in the respective papers implementing intent-awareness in a system; we will discuss these technical approaches in more detail in Section 3. Following the spirit of early diversity-oriented IARS approaches, the authors of (Kaya and Bridge, 2019a) for example state that “[IARS] ensure that the set of recommendations contains items that cover each of the user’s interests […].” More recent works like (Wang et al., 2020c) relate intents with observed actions and provide examples, in this case for shopping intents in the e-commerce domain, e.g., “passing time” or “shopping for others”. In several other works, e.g., (Chen et al., 2020, 2019), the term ‘intent’ is mostly considered to be equivalent with (observed) interest, e.g., in specific topics or item categories. Finally, in some works like (Jannach et al., 2017), the term (short-term) ‘intent’ is merely used to describe a sequential recommendation approach that puts more emphasis on the last observed user interactions. Thus, intent is sometimes solely equated with observed user actions, without additional consideration of users’ underlying motivations for these actions.
@@
Given the examples from the previous paragraph, we can roughly categorize existing IARS approaches into three classes. First, a comparably small number of the surveyed works are based on a small set of pre-defined and application-specific explicit user intents, e.g., “discover new music to listen right now” or “find background music for listen to” (Mehrotra et al., 2019). In such an approach, one can for example apply different recommendation strategies depending on the currently assumed intent. Second, some other works provide examples of such explicit intents for illustration, e.g., “shopping for others” (Wang et al., 2020c), but are actually based on the concept of ‘latent’ intents. In fact, much of the surveyed literature is based on the idea of incorporating (additional) latent variables in the models to be able to account for a typically small set of possible user intents. Third, there are a number of approaches, as mentioned above, in which intent is considered to be roughly equivalent to observed past interests in categories, certain item features, or even individual items. In such approaches, additional variables are commonly introduced, which are however not assumed to directly correspond to specific underlying intents in the sense of goals a user wants to achieve. Instead, such variables may, for instance, model the user’s time-varying interest in certain item categories or other assumed relationships between individual items. We denote such approaches implicit intent models.
@@
An IARS is a recommender system that is designed to capture the users’ underlying current motivations and goals in order to support them.
Our teleological definition is purposely rather broad, yet it expresses that the design of any IARS should be driven by the ambition to support the users’ current motivations. We emphasize that we use the term ‘intent’ to exclusively refer to the user’s immediate or short-term motivations. Users may certainly have diverse preferences and needs over time. However, the core goal of an IARS is to identify the current need or intent when a user interacts with the recommendation service.
Given the breadth of different notions of intent-awareness mentioned above, a multitude of technical ways exist to implement intent-awareness. On one end of the spectrum, there are approaches based on an explicit list of pre-defined application-specific intents. In contrast, there are also works in which a sequential recommender system that puts more emphasis on the most recently observed interactions is seen as a mechanism to implement an (implicitly) intent-aware system, e.g. (Jannach et al., 2017).
One might argue that any sequential next-item prediction model may be thought of as being intent-aware. We however recall that our definition is based on the design goals of a model. Let us consider Amazon’s “Customers who bought …also bought” approach as an example for a basic next-item recommendation technique. The recommendations returned for a given item by this purely statistical approach may either be alternatives (similar movies on a video-on-demand service) or complementary items (accessories for a smartphone on an e-commerce shop). In our interpretation, such an approach would therefore not be considered intent-aware. The reason is that there is no element in the underlying design of the approach that aims at capturing the user’s current motivations or intent, i.e., whether the user looks for alternatives or complementary items.
@@ 2.2. Research Methodology
We adopted a semi-systematic approach to identify relevant papers. First, we queried services like Google Scholar and digital libraries for scientific papers that mention the terms ‘intent’ or ‘intention’ together with the term ‘recommend’ in the title or abstracts. We then screened these papers for relevant ones and followed the references mentioned in these papers to identify additional papers. We only retained papers that had undergone a peer-review process, i.e., we excluded preprints from our analysis. Furthermore, a paper was considered relevant if its contribution was motivated by the idea of intent-awareness in the sense of our definition from above even though it did not explicitly contain our search terms in the title or abstract. As such, a small number of papers was considered that emphasized on identifying and addressing the user’s current aims and interests. We also identified a number of works which mention the term ‘intent’ in the title or abstract, but not in a context that is relevant for our survey. For example, some works aim at predicting if a consumer will make a purchase in an ongoing session or not, e.g., (Requena et al., 2020). In such works, the goal is not to make recommendations that match a certain intent, as done in IARS, but to assess the general propensity of a consumer to make a purchase, see also (Koniew, 2020).
Overall, we ended up with 85 papers that we considered relevant and which we categorize later in Section 3. A possible limitation of our work is that our survey is not based on the strict procedures of a systematic review in the sense of Kitchenham et al. (Kitchenham et al., 2009; Kitchenham and Brereton, 2013). Nonetheless, by using relevant queries and by applying a snowballing procedure to find additional papers, we are confident that our survey comprises a representative collection of papers reflecting the main areas of IARS research and capturing the conceptual breadth of the field. Our literature search also surfaced surveys that did not propose a technical contribution or reported an experimental study. We discuss these papers separately next.
@@
Surveys on related the areas, as discussed above, such as time-aware, contextual, and sequential recommender systems, can be found in (Campos et al., 2014; Bogina et al., 2022; Adomavicius and Tuzhilin, 2015; Kaminskas and Bridge, 2016; Quadrana et al., 2018; Wang et al., 2021a). Early intent-aware recommendation systems were inspired by existing approaches to search result diversification in the field of Information Retrieval, see (Santos et al., 2015) for a related survey. Indeed, considering the users’ intents, e.g., in terms of their ‘informational needs’ has a long history in Information Retrieval, see (Broder, 2002) for an early taxonomy. A review of intent-awareness in the particular context of multimedia information retrieval can be found in (Kofler et al., 2017). Like in our survey, Kofler at al. (Kofler et al., 2017) found that various notions of intent-awareness exist in the literature. Our definition of intent-awareness is similar to theirs by focusing on the ‘why’ dimension, i.e., the reason and goal behind a search, and not the ‘what’ dimension of information needs. In the area of music information retrieval, Schedl et al. (Schedl et al., 2018) reviewed current challenges and visions in music recommendation research. While their review work is not primarily focused on the topic of user intents, they identify an improved understanding of listening intents and the user’s purpose as a grand challenge in this area. In that context, they also mention that creators may have specific intents in mind when designing a playlist (e.g., for relaxing).
Probably most similar to our survey is the work by Rafique et al. (Rafique et al., 2023), which also focuses on intent-aware recommender systems. They, however, study the current and future role of IARS from the lens of smart cities. A particular emphasis is put on the role of recent technological developments in this context, e.g., the increasing usage of wearable and intelligent devices, the Internet of Things or Edge Computing approaches. As a result, new opportunities arise to build future intent-aware systems based, for instance, on real-time activity, emotion tracking or the explicit elicitation of user intents based on new sensors and devices. In their survey, the authors thus particularly focus on case studies of IARS in smart cities and on key requirements in this context.
@@ 3. Technical Approaches for IARS
In this section, we categorize the considered studies on intent-aware recommender systems according to their distinguishing algorithmic properties. Early Profile Diversity Matching approaches have a particular optimization goal and rely on item-metadata. Sequence- and Time-Aware models are based on the availability of temporal information about past user interactions. Finally, Latent Intent and Explicit Intent Modeling approaches assume the existence of a small set of common intents in a given application use case. Our resulting categorization is illustrated in Figure 1.
@@
We note that some papers may fall into more than one category, e.g., a latent-intent modeling approach with the goal of diversification. In such cases, we categorized these works according to what we identified as their main conceptual contribution.